OpenStudy (anonymous):

An event lasts 1,560,000 milliseconds. Approximately how long did the event last in hours

OpenStudy (anonymous):

First note that using dimensional analysis: \[(\frac{ms}{1})(\frac{1000s}{ms})(\frac{1m}{60s})(\frac{1hr}{60m})\] Will Convert a unit from ms to hours. Notice how all units cancell each other out except the hours.

OpenStudy (anonymous):

Hold on, typo

OpenStudy (anonymous):

\[(\frac{ms}{1})(\frac{1s}{1000ms})(\frac{1m}{60s})(\frac{1hr}{60m})\]

OpenStudy (anonymous):

Thank you!(: